# Late-blooming magnetars: awakening as long period transients after a dormant cooling epoch

### v2

@@ A.1 Crustal failures
Perhaps more importantly, Bransgrove, Levin & Beloborodov (2025) note that by tracking Maxwell stresses (as in Sec. III), rather than mechanical ones directly, overestimates for event rates may be obtained. While modifying the code to compute Lagrangian displacements to more realistically estimate stresses lies beyond our current capabilities, we provide here a worked example that illustrates for mixed poloidal-toroidal fields of comparable strength the rate estimates are relatively insensitive to the exact criterion used.
As discussed by Bransgrove, Levin & Beloborodov (2018) and others, the strong radial stratification in the crust permits only horizontal ($\xi_{r}=0$) and incompressible ($\nabla\cdot\boldsymbol{\xi}=0$) deformations because the hydrostatic pressure is much larger than the Coulomb energy density of the crystal. In the axisymmetric limit, the Lagrangian deformation thus reads
@@
Following Pons & Geppert (2007) we write
@@
which shows effectively that the stress is sourced by poloidal-toroidal cross-terms, while the Maxwell stress is instead sourced by magnetic pressures (which could indeed be much larger). Implementing a routine that mimics that described in Sec. III for resetting the crust following an overstraining, we compare three cases here: (i) $|\mathcal{M}_{ij}-\mathcal{M}_{ij}^{\rm eq}|$ for the pure Maxwell stress (as in Pons & Perna, 2011; Perna & Pons, 2011; Dehman et al., 2020), (ii) $|\mathcal{M}^{\rm T}_{ij}-\mathcal{M}^{\rm T,eq}_{ij}|$ for the traceless component of the Maxwell stress (an adjustment we have made in this paper; see equation 3), or (iii) $|\boldsymbol{\tau}_{ij}-\boldsymbol{\tau}^{\rm eq}_{ij}|$ for the true mechanical strain (A3) determined by solving Eq. (A4).
Table 1: Relevant details for a variety of (plausibly) isolated LPTs (cf. Suvorov, Dehman & Pons, 2025), namely upper limits on the X-ray luminosity ($L_{\rm X}^{\rm max}$), the observed pulse period ($P$; in descending order) and its (upper-limit) derivative ($\dot{P}^{\rm max}$), together with estimated duty cycles ($D$) and distances ($d$). Additional details on DA J1832 are provided in Tab. 2. Some object names have been shortened.
@@
For the sake of demonstration, we choose simple flux profiles that resemble that from Qu & Bransgrove (2026). Explicitly, suppose $\Phi_{1}=B_{p}\left(a_{2}r^{2}+a_{4}r^{4}\right)$ and $\Psi_{1}=B_{t}\left(b_{2}r^{2}+b_{4}r^{4}\right)$ where the constants $a_{i}$ are chosen to match to a force-free dipole at the surface and $b_{i}$ are chosen to make the toroidal field vanish there. The prefactors $B_{p}$ and $B_{t}$ set the strengths of the poloidal and toroidal components, respectively, which we choose to abide by the phenomenological volume-averaged solution to the induction equation introduced by Aguilera, Pons & Miralles (2008),
@@
For transparency, we close this section by noting that there are some important caveats worth highlighting with respect to our models. Probably the most critical aspect is that there is no explicit inclusion of elasticity or plasticity: failures are handled through an instantaneous, von Mises mechanism where the crust heals prior to the next time-step without undergoing any lateral motions. In particular, while we have used the von Mises criterion to determine failure sites, molecular dynamics simulations instead indicate that a more physical mechanism is the Zhurkov model (Chugunov & Horowitz, 2010), where failures occur when thermodynamic fluctuations exceed a threshold energy over finite time intervals. This leads to gradual deformations rather than immediate failures and is more appropriate for modelling elastic/plastic transitions, though is difficult to handle numerically owing to the timescales of the problem. Having the crust reset instantly is also not really consistent with the radio-activation picture, since the zone should remain plastic for multiple time steps (and hence not accumulate stress) to account for a non-negligible duty cycle (Cooper & Wadiasingh, 2024). Straining motions should also contribute a Faraday-like $\boldsymbol{v}_{\rm ion}\times\boldsymbol{B}$ term within equation (1) (Lander & Gourgouliatos, 2019; Bransgrove, Beloborodov & Levin, 2020), though we do not explicitly account for this feature in this first magnetothermal study of isolated LPTs.
@@ Appendix B Relevant properties of isolated LPTs
GPM J1839 has tight timing constraints (Men et al., 2025) ($\dot{P}\lesssim 3.6\times 10^{-13}\text{ ss}^{-1}$), spanning the last $\sim 3$ decades, set while the source was intermittently radio-loud. Owing to the mild limits on this source’s X-ray luminosity (Men et al., 2025), shallow failures triggering radio activity (Cooper & Wadiasingh, 2024) remain plausible provided the present-day field is comparatively weak ($\lesssim 10^{14}$ G) and that the bulk of the enhanced spindown occurred prior to the present epoch when the magnetoelastic energy reservoir was plentiful.
@@ B.1 DA J1832 activity timeline and alternative scenarios
Note in particular that X-ray data were unavailable during the onset of radio pulsing in December 2023. Although one could take the observations at face value (see below), since the absence of evidence is not evidence of absence we have instead hypothesised the following: a (sub)-crustal event took place around the start of December 2023, which was directly responsible for radio activation by injecting magnetospheric twist or initiating a plastic flow. Thermoplastic heating or particle backflow (Beloborodov & Levin, 2014) generated a localised hotspot surrounding the failure site as the twist ramped before decaying in the $\lesssim$ months following February 2024. A schematic illustrating this scenario is depicted in Fig. 7, while a summary of relevant source parameters are listed in Tab. 2.
@@ Appendix C Burst-to-spindown efficiencies
where $\zeta_{1}$ represents an intermediate conversion factor between (magneto-)elastic and outburst energies, with $\zeta_{2}$ then being the fraction of outburst energy that contributes to spindown. Note that our $\bar{\zeta}$ is phenomenologically similar to that used in other studies of enhanced spindown (see, e.g., figure 1 in Beniamini, Wadiasingh & Metzger, 2020).
The global simulations by Bransgrove, Beloborodov & Levin (2020) of pulsar quakes where the crust and core are coupled – as appropriate to CT models – suggest that $\sim 1\%$ of the quake energy escapes as observable radiation (i.e., $\zeta_{1}\sim 10^{-2}$). This radiation escapes as waves which may not directly bolster the dipole moment, but gives us a handle on how to relate numerically-determined quake energy to observable radiation. For magnetars, $\zeta_{1}$ may be higher as the ambient magnetic field mitigates thermodynamic losses: Blaes et al. (1989) calculated the crust-to-magnetosphere transmission coefficient, finding it scales with the field strength at high frequencies, especially in cases where the affected region liquefies. It could reach $\sim\mathcal{O}(10\%)$ for $B\gtrsim 10^{14}$ G.

### v3

@@ A.1 Crustal failures
Perhaps more importantly, Bransgrove et al. (2025) note that by tracking Maxwell stresses (as in Sec. III), rather than mechanical ones directly, overestimates for event rates may be obtained. While modifying the code to compute Lagrangian displacements to more realistically estimate stresses lies beyond our current capabilities, we provide here a worked example that illustrates for mixed poloidal-toroidal fields of comparable strength the rate estimates are relatively insensitive to the exact criterion used.
As discussed by Bransgrove et al. (2018) and others, the strong radial stratification in the crust permits only horizontal ($\xi_{r}=0$) and incompressible ($\nabla\cdot\boldsymbol{\xi}=0$) deformations because the hydrostatic pressure is much larger than the Coulomb energy density of the crystal. In the axisymmetric limit, the Lagrangian deformation thus reads
@@
Following Pons and Geppert (2007) we write
@@
which shows effectively that the stress is sourced by poloidal-toroidal cross-terms, while the Maxwell stress is instead sourced by magnetic pressures (which could indeed be much larger). Implementing a routine that mimics that described in Sec. III for resetting the crust following an overstraining, we compare three cases here: (i) $|\mathcal{M}_{ij}-\mathcal{M}_{ij}^{\rm eq}|$ for the pure Maxwell stress (as in Pons and Perna, 2011; Perna and Pons, 2011; Dehman et al., 2020), (ii) $|\mathcal{M}^{\rm T}_{ij}-\mathcal{M}^{\rm T,eq}_{ij}|$ for the traceless component of the Maxwell stress (an adjustment we have made in this paper; see equation 3), or (iii) $|\boldsymbol{\tau}_{ij}-\boldsymbol{\tau}^{\rm eq}_{ij}|$ for the true mechanical strain (A3) determined by solving Eq. (A4).
Table 1: Relevant details for a variety of (plausibly) isolated LPTs (cf. Suvorov et al., 2025), namely upper limits on the X-ray luminosity ($L_{\rm X}^{\rm max}$), the observed pulse period ($P$; in descending order) and its (upper-limit) derivative ($\dot{P}^{\rm max}$), together with estimated duty cycles ($D$) and distances ($d$). Additional details on DA J1832 are provided in Tab. 2. Some object names have been shortened.
@@
For the sake of demonstration, we choose simple flux profiles that resemble that from Qu and Bransgrove (2025). Explicitly, suppose $\Phi_{1}=B_{p}\left(a_{2}r^{2}+a_{4}r^{4}\right)$ and $\Psi_{1}=B_{t}\left(b_{2}r^{2}+b_{4}r^{4}\right)$ where the constants $a_{i}$ are chosen to match to a force-free dipole at the surface and $b_{i}$ are chosen to make the toroidal field vanish there. The prefactors $B_{p}$ and $B_{t}$ set the strengths of the poloidal and toroidal components, respectively, which we choose to abide by the phenomenological volume-averaged solution to the induction equation introduced by Aguilera et al. (2008),
@@
For transparency, we close this section by noting that there are some important caveats worth highlighting with respect to our models. Probably the most critical aspect is that there is no explicit inclusion of elasticity or plasticity: failures are handled through an instantaneous, von Mises mechanism where the crust heals prior to the next time-step without undergoing any lateral motions. In particular, while we have used the von Mises criterion to determine failure sites, molecular dynamics simulations instead indicate that a more physical mechanism is the Zhurkov model (Chugunov and Horowitz, 2010), where failures occur when thermodynamic fluctuations exceed a threshold energy over finite time intervals. This leads to gradual deformations rather than immediate failures and is more appropriate for modelling elastic/plastic transitions, though is difficult to handle numerically owing to the timescales of the problem. Having the crust reset instantly is also not really consistent with the radio-activation picture, since the zone should remain plastic for multiple time steps (and hence not accumulate stress) to account for a non-negligible duty cycle (Cooper and Wadiasingh, 2024). Straining motions should also contribute a Faraday-like $\boldsymbol{v}_{\rm ion}\times\boldsymbol{B}$ term within equation (1) (Lander and Gourgouliatos, 2019; Bransgrove et al., 2020), though we do not explicitly account for this feature in this first magnetothermal study of isolated LPTs.
@@ Appendix B Relevant properties of isolated LPTs
GPM J1839 has tight timing constraints (Men et al., 2025) ($\dot{P}\lesssim 3.6\times 10^{-13}\text{ ss}^{-1}$), spanning the last $\sim 3$ decades, set while the source was intermittently radio-loud. Owing to the mild limits on this source’s X-ray luminosity (Men et al., 2025), shallow failures triggering radio activity (Cooper and Wadiasingh, 2024) remain plausible provided the present-day field is comparatively weak ($\lesssim 10^{14}$ G) and that the bulk of the enhanced spindown occurred prior to the present epoch when the magnetoelastic energy reservoir was plentiful.
@@ B.1 DA J1832 activity timeline and alternative scenarios
Note in particular that X-ray data were unavailable during the onset of radio pulsing in December 2023. Although one could take the observations at face value (see below), since the absence of evidence is not evidence of absence we have instead hypothesised the following: a (sub)-crustal event took place around the start of December 2023, which was directly responsible for radio activation by injecting magnetospheric twist or initiating a plastic flow. Thermoplastic heating or particle backflow (Beloborodov and Levin, 2014) generated a localised hotspot surrounding the failure site as the twist ramped before decaying in the $\lesssim$ months following February 2024. A schematic illustrating this scenario is depicted in Fig. 7, while a summary of relevant source parameters are listed in Tab. 2.
@@ Appendix C Burst-to-spindown efficiencies
where $\zeta_{1}$ represents an intermediate conversion factor between (magneto-)elastic and outburst energies, with $\zeta_{2}$ then being the fraction of outburst energy that contributes to spindown. Note that our $\bar{\zeta}$ is phenomenologically similar to that used in other studies of enhanced spindown (see, e.g., figure 1 in Beniamini et al., 2020).
The global simulations by Bransgrove et al. (2020) of pulsar quakes where the crust and core are coupled – as appropriate to CT models – suggest that $\sim 1\%$ of the quake energy escapes as observable radiation (i.e., $\zeta_{1}\sim 10^{-2}$). This radiation escapes as waves which may not directly bolster the dipole moment, but gives us a handle on how to relate numerically-determined quake energy to observable radiation. For magnetars, $\zeta_{1}$ may be higher as the ambient magnetic field mitigates thermodynamic losses: Blaes et al. (1989) calculated the crust-to-magnetosphere transmission coefficient, finding it scales with the field strength at high frequencies, especially in cases where the affected region liquefies. It could reach $\sim\mathcal{O}(10\%)$ for $B\gtrsim 10^{14}$ G.